# Cone Ranking for Multi-Criteria Decision Makingthanks: Paper accepted and presented by D. Kostner at MODeM Workshop at ECAI 2023, Krakow, Poland

### 04 Dec 2023 v1

@@ 6 Supervised classification models
$n\in\mathrm{I\negthinspace N}$, for this purpose. These level sets correspond to the set-valued cone quantiles introduced in [6]. One can ask which is the greatest number $n$ satisfying #$\#\{x\in X\mid x\in L_{X,C}(n)\}\geq\frac{N\alpha}{100}$ which even gives some information about the set $X$ of alternatives as a whole: the higher this number is (in relation to the number of alternatives), the bigger is the group of ”good” alternatives.
@@
A related questions is to cluster the alternatives in the ”really good ones,” ”the really bad ones” and ”the ugly ones” where the latter category is meant to include those alternatives which considerably deviate from the rest in a good way w.r.t. some of the criteria and in a bad way w.r.t. other criteria. For this purpose, one can consider the sets $L_{X,C}(N/2)$ (the good ones, i.e., with a ranking of at least $N/2$), $L_{X,-C}(N/2)$ (the bad ones since the direction of the cone is reverted) and $X\backslash(L_{X,C}(N/2)\cup L_{X,-C}(N/2))$. The set $L_{X,C}(N/2)$ corresponds to the multivariate median from [6]. It should be noted that the sets $L_{X,C}(N/2)$, $L_{X,-C}(N/2)$ could even be empty (examples can be found in [7]).
@@
Again, the shape of these sets also gives information on the set of alternatives as a whole and can be utilized, for example, for machine learning procedures such as supervised classification models as defined in [8]. Such models take the cone $C$ as an input and adapt the parameter $n$ in order to find the set $L_{X,C}(n)$ that achieves the lowest error rate based on the labeled data. This machine learning concept can be used as a recommender system: the alternatives in $L_{X,C}(n)$ will be recommended where the cone $C$ stems from the given preference relation of the customer/user. The goal is to find the set $L_{X,C}(n)$, by adjusting $n$, that includes the alternatives that will be accepted (with a high probability). Therefore, labeled data are necessary. In this case, the alternatives which are represented by their attributes, get a label that provides the information whether or not they are acceptable. The further development into a semi-supervised model is straightforward as the unlabeled data can be assigned to the labeled one via the order relation $\leq_{C}$. The goal is to determine the set $L_{X,C}(n)$ and the value $n$, respectively, that minimizes the classification error rate.
@@
If there are unlabeled alternatives dominating already positively labeled ones w. r. t. $\leq_{C}$, then these are also labeled as acceptable. The same principle is applied in the opposite direction for undesirable data. If the amount of unlabeled data is smaller then the labeled one and some of the former data is not comparable with the latter one, then this unlabeled data can be omitted for the calculation of the set $L_{X,C}(n)$ and reinserted as new data later on. The challenge is to create an algorithm that determines the set $L_{X,C}(n)$ and the value $n$, respectively, that minimizes the classification error rate. The false positives as well as false negatives should be minimized. This could be achieved by adjusting $n$ via appropriate methods. The resulting set $L_{X,C}(n)$ is then used to classify new data points. Of course, this set can be updated after a specified threshold of new data.
@@ 7 Conclusion and perspective
A new ranking function is proposed which is derived from a statistical function [6] which in turn resembles and generalizes the so-called half-space or Tukey depth function. In statistics, a variety of more recently introduced depth functions with different properties exist. The ranking function proposed in Definition 2.1 is based on a cone version of the half-space depth function. The same idea can be applied to other statistical depth functions like the zonoid depth [13] or expectile depth functions [2]. This may lead to a variety of new ranking functions for MCDM. The computation of the values of $r_{X,C}$ is a non-trivial task but can be done based on a merge of sorting algorithms with convex geometry methods, compare [7] for first impressions and remarks on complexity. The proposed ranking methods give a ranking of a given set of alternatives relative to each other. Therefore, rank reversal features appear naturally, but can be characterized and used in order to analyse the decision making procedure itself.

### 14 Jan 2024 v2

@@ 6 Classification models
$n\in\mathrm{I\negthinspace N}$, for this purpose. These level sets correspond to the set-valued cone quantiles introduced in [10]. One can ask which is the greatest number $n$ satisfying #$\#\{x\in X\mid x\in L_{X,C}(n)\}\geq\frac{N\alpha}{100}$ which even gives some information about the set $X$ of alternatives as a whole: the higher this number is (in relation to the number of alternatives), the bigger is the group of “good” alternatives.
@@
Unsupervised classification. A related questions is how to cluster the alternatives in the “really good ones,” “the really bad ones” and “the ugly ones” where the latter category is meant to include those alternatives which considerably deviate from the rest in a good way w.r.t. some of the criteria and in a bad way w.r.t. other criteria. For this purpose, one can consider the sets $L_{X,C}(N/2)$ (the good ones, i.e., with a ranking of at least $N/2$), $L_{X,-C}(N/2)$ (the bad ones since the direction of the cone is reverted) and $X\backslash(L_{X,C}(N/2)\cup L_{X,-C}(N/2))$. The set $L_{X,C}(N/2)$ corresponds to the multivariate median from [10]. It should be noted that the sets $L_{X,C}(N/2)$, $L_{X,-C}(N/2)$ could even be empty (examples can be found in [11]).
Again, the shape of these sets also gives information on the set of alternatives as a whole and can be utilized, for example, for machine learning procedures such as unsupervised classification models as defined in [8]. Such models take the cone $C$ and the parameter $n$ as an input in order to partition the data. In this case, the cone $C$ can be seen as an externally given ordering cone reflecting the preference of a decision maker and $n$ defines how strict this preference is applied to the data. The output of this method are the three labels discussed above: good, bad, and ugly, the latter one a new concept reflecting the presence of non-comparable alternatives.
Supervised classification. Reversing the question, a supervised classification model can be developed by finding the set $L_{X,C}(n)$ that achieves the lowest error rate if the data $X$ is labeled and the cone $C$, the parameter $n$, or both are not given as an input.
@@
The following recommender system can serve as blueprint where the cone $C$ is given and the parameter $n$ must be determined: the alternatives in $L_{X,C}(n)$ will be recommended where the cone $C$ stems from the given preference relation of the customer/user. The goal is to find, by adjusting $n$, the set $L_{X,C}(n)$ that includes the alternatives that will be accepted (with a high probability). In this case, the alternatives which are represented by their attributes, get a label that provides the information whether or not they are acceptable. The goal is to determine the set $L_{X,C}(n)$ and the value $n$, respectively, that minimizes the classification error rate. By comparing this implementation of $L_{X,C}(n)$ with the logistic regression, as described in [14], two observations can be made. Firstly, the logistic regression has a response variable/label $Y$ and a predictor $X$ as input, whereas the new method has the preference information $C$ as additional input. The preference information can be seen as an additional tuning/customization of the model. Secondly, the parameter $n$ can also be seen as a threshold (probability cut-off value) as it affects the classification into the labels.
@@
The further development into a semi-supervised model is straightforward as the unlabeled data can be assigned to the labeled one via the order relation $\leq_{C}$. If there are unlabeled alternatives dominating already positively labeled ones w.r.t. $\leq_{C}$, then these are also labeled as acceptable. The same principle is applied in the opposite direction for undesirable data. If the amount of unlabeled data is smaller then the labeled one and some of the former data is not comparable with the latter one, then this unlabeled data can be omitted for the calculation of the set $L_{X,C}(n)$ and reinserted as new data later on. The challenge is to create an algorithm that determines the set $L_{X,C}(n)$ and the value $n$, respectively, that minimizes the classification error rate. The false positives as well as false negatives should be minimized. This could be achieved by adjusting $n$ via appropriate methods. The resulting set $L_{X,C}(n)$ is then used to classify new data points. Of course, this set can be updated after a specified threshold of new data.
Towards preference learning. If the cone $C$ is understood as a model for a customer preference, one might want to learn about it [6]. If there are two classes inside $X$, the parameter $n$ can be extrapolated from the number of elements in each class. This simple method works especially well if the cone $C$ is “aligned” with the data $X$ producing level sets with low classification error rates: compare Figure 6.1. However, if the cone $C$, seen as a first estimate of the customer preference, is not aligned with the data $X$, the level sets could not include all relevant alternatives resulting in poor classification, see Figure 6.3. This could be resolved by rotating the cone $C$ towards the data $X$.
Figure 6.1: $X$ and $C$ are aligned.
The next idea not only looks for an optimal parameter $n$, but also partially adjusts the cone $C$ to the data. The linear Support Vector Machine (SVM) is used to align the cone $C$ with the data. The SVM separates the classes with a maximal margin hyperplane that. The idea is to adjust the cone $C$ to the data $X$ by rotating it towards the normal vector of the maximal margin hyperplane $w_{SVM}$ as in Figure 6.3.
Figure 6.2: $X$ and $C$ are not aligned.
Figure 6.2: $X$ and $C$ are not aligned.
Figure 6.3: Alignment of $C$ via SVM
In higher dimensions ($d>2$) this can be an elaborated procedure. For the computation of $L_{X,C}(n)$ the vertex representation of the dual cone $C^{+}$ is needed. This can be used to find an interior direction $w_{int}$ of $C^{+}$ by taking the convex combination of the extreme directions of $C^{+}$. Now, the rotation matrix from $w_{int}$ to $w_{SVM}$ is calculated. This matrix is used to rotate the dual cone $C^{+}$ towards $w_{SVM}$. This method only works if the two classes can be clearly separated by a hyperplane. This process can also be interpreted as a refinement of a classification via the SVM by introducing the order (preference) given by the “data adjusted” cone $C$.
Of course, the cone $C$ can be adapted to the data not only via a rotation, but also by changing its general shape. However, already in three dimensions, there are infinitely many degrees of freedom to change the cone. We consider this question a very interesting challenge since it gives a new direction of research linking the analysis of “mere” data with preference relations. It brings the statistical concepts from [10], especially the level sets $L_{X,C}(n)$, to the field of preference learning as presented, e.g., in LABEL:FurnkranzHullermeier10.
@@ 7 Conclusion and perspective
A new ranking function is proposed which is derived from a statistical function [10] which in turn resembles and generalizes the so-called half-space or Tukey depth function. In statistics, a variety of more recently introduced depth functions with different properties exist. The ranking function proposed in Definition 2.1 is based on a cone version of the half-space depth function. The same idea can be applied to other statistical depth functions like the zonoid depth [19] or expectile depth functions [12]. This may lead to a variety of new ranking functions for MCDM. The computation of the values of $r_{X,C}$ is a non-trivial task but can be done based on a merge of sorting algorithms with convex geometry methods, compare [11] for first impressions and remarks on complexity. The proposed ranking methods give a ranking of a given set of alternatives relative to each other. Therefore, rank reversal features appear naturally, but can be characterized and used in order to analyse the decision making procedure itself. Extensions to set preferences open the path to completely new comparison methods, e.g., for evolutionary multiobjective optimization on the one hand and new scalarization methods for multi-criteria optimization problems on the other hand. Data analysis for ordered data via statistical learning procedures are within reach based on the concepts in this paper.